# BWCache: Accelerating Video Diffusion Transformers through Block-Wise Caching

### v2

@@ 3.4 Periodic Cache Recomputation
Continuously reusing the same block may result in latent drift and gradually erase fine-grained details Liu et al. (2025b). To prevent this cache stagnation, we adopt the progressive computation strategy from PAB Zhao et al. (2024a): within the caching interval, each DiT block is periodically recomputed at a defined reuse interval . Specifically, after computing the block at time step , its features are cached and reused for the subsequent steps (i.e., ), and the block is then updated at step . This process is formalized as:
@@ 4.1 Experiment Setup
Base Models and Compared Methods. We apply the proposed acceleration techniques to various video generation diffusion models, including Open-Sora Zheng et al. (2024), Open-Sora-Plan Lin et al. (2024), and Latte Ma et al. (2025). We compare our method with recent effective video generation techniques such as PAB Zhao et al. (2024a), T-GATE Zhang et al. (2024b), -DiT Chen et al. (2024b), and TeaCache Liu et al. (2024) to highlight our advantages. The inference config of base models in Appendix 1.
Evaluation Metrics and Datasets. To evaluate the performance of video generation acceleration methods, we primarily focus on two aspects: visual quality and inference efficiency. We utilize VBench Huang et al. (2024b), LPIPS Zhang et al. (2018), PSNR, and SSIM Wang and Bovik (2002) for visual quality assessment. VBench is a comprehensive benchmarking suite for video generation models. LPIPS, PSNR, and SSIM measure the similarity between videos generated by the accelerated methods and the original models. Several evaluation metrics are detailed in Appendix 2. To assess inference efficiency, we use Floating Point Operations (FLOPs) and inference latency as metrics. In our main experiments, we generate over 1,000 videos for each model under each baseline using prompts sourced from Open-Sora gallery Lin et al. (2024), VBench benchmark Huang et al. (2024b), and T2V-CompBench Sun et al. (2024).
@@ 4.2 Main Results
Quantitative Comparison. The experimental results in Table 1 demonstrate the effectiveness of BWCache across multiple DiT-based video generation models. Our proposed BWCache outperforms existing optimization methods in both visual quality and efficiency. For instance, on Open-Sora, BWCache achieves advanced visual metrics while reducing latency by 1.61 compared to the original model. Notably, BWCache surpasses the training-free TeaCache method and maintains competitive acceleration performance. The block-wise method of BWCache leverages DiT block redundancies more effectively without sacrificing visual details. The timestep-level caching of TeaCache is less adaptive to intra-timestep variations, leading to higher quality degradation. It is worth noting that TeaCache achieves faster acceleration than BWCache on the Open-Sora-Plan model. This is because there is significant variation in block features during the early inference stages of the Open-Sora-Plan model, and TeaCache focuses solely on input features while neglecting this variation. This aggressive acceleration strategy results in its generation quality lagging far behind that of BWCache. More experiments can be found in Appendix 3.
Visual Quality Comparison. As shown in Figure 4, we visualize the videos generated by our method compared to the original model, PAB, and TeaCache. The results demonstrate that BWCache outperforms other methods in visual quality with lower latency. More visual results are illustrated in Appendix 4.
@@ 4.3 Scaling Abilities
Scaling to Multiple GPUs. Table 2 summarizes the inference efficiency of various methods when scaled across multiple GPUs using Dynamic Sequence Parallelism (DSP) Zhao et al. (2024b). The experiments, conducted using several models, reveal that BWCache consistently outperforms both the original model and PAB in terms of latency reduction across all tested scenarios. Specifically, BWCache achieves the lowest latencies, demonstrating significant speed-ups as the number of GPUs increases. These results validate BWCache as a universally effective acceleration method for DiT-based models under multi-GPU deployments.
@@
| Ours | 14.4(1.8) | 7.3(3.6) | 4.5(5.8) | 4.0(6.5) |
Performance at Different Length and Resolution. Figure 5 demonstrates the inference efficiency of BWCache across various video lengths and resolutions using the Open-Sora model. The results show that latency increases significantly as the length and resolution of the video increased, but BWCache consistently performs better than the original model in all configurations. BWCache exhibits particularly notable acceleration advantages in processing high-resolution long videos. For example, with eight GPUs, BWCache attains a latency of just 11.08 seconds for the Open-Sora model with 204 frames at 480P, representing a remarkable 17.16 speed-up compared to the base model. This trend is also reflected in other configurations, where BWCache consistently provides substantial improvements in efficiency. Overall, these results validate its capability to improve inference speeds across various video lengths and resolutions, making it an efficient solution for real-time video generation.
@@ 4.4 Ablation Studies
Quality-Efficiency Trade-Off. Figure 1 compares the quality-latency trade-off of BWCache, PAB, and TeaCache. The indicator threshold in Eq.(7) is 0.15, 0.20, and 0.25. The results clearly demonstrate that BWCache outperforms PAB and TeaCache in visual quality across all evaluated metrics, indicating a notable improvement in the fidelity of video generation. For each metric, BWCache not only achieves higher quality scores but also does so while maintaining lower latency, which reflects its efficient processing capabilities. Overall, BWCache achieves improved visual quality at lower latency than PAB and TeaCache.
@@
Reuse Rate. Table 3 illustrates how varying the reuse rates affects performance. It reveals that the “Fast" configuration achieves the highest reuse rate at 59.00%, resulting in the lowest latency of 19.16 seconds. The “Slow" configuration only had a reuse rate of 41.38%, but the visual quality is the highest among all configurations.
Visualized results in Figure 6 complement these findings by providing the reuse rates of blocks across each timestep for various thresholds. This figure illustrates the dynamic nature of block reuse at different timesteps, with evident fluctuations that suggest varying block reuse efficiency at different stages of video generation. In general, the block features in the middle stage of the generation change less, and a high reuse rate is obtained, which is consistent with our previous analysis.
@@ 5 Conclusion
In this paper, we introduced BWCache, a novel caching method designed to accelerate video Diffusion Transformers. By analyzing feature dynamics in Spatial-Temporal DiT blocks, we established that block features across timesteps had significant computational redundancy. BWCache addressed this by employing dynamic block-wise caching and a similarity indicator, which selectively reused block features when their differences fell below a predefined threshold. Periodic recomputation was employed to mitigate potential latent drift. Extensive experiments demonstrated that BWCache achieved robust efficiency and visual quality across diverse generation models, sampling schedules, and video parameters. These results highlighted its potential for real-world applications. The current indicator threshold was preset, and future work will dynamically adjust the threshold according to different generation tasks.
Appendix A Experimental Settings
Base Model. This paper focuses on DiT-based video generation, evaluating three state-of-the-art open-source models: Open-Sora Zheng et al. [2024], Open-Sora-Plan Lin et al. [2024], and Latte Ma et al. [2025]. Open-Sora employs Spatial-Temporal DiT Ho et al. [2022], allowing for the creation of videos up to 16 seconds long at 720p resolution. Open-Sora Plan focuses on generating high-resolution videos with extended durations based on diverse user inputs, featuring components like a Wavelet-Flow Variational Autoencoder and specialized denoisers. Latte effectively extracts spatio-temporal tokens from videos and utilizes DiT blocks to model the token distribution in a latent space. The inference configs of three models are shown in Table 5, which strictly follow the official settings.
@@ Appendix A Experimental Settings
Appendix B Metrics
@@
Learned Perceptual Image Patch Similarity (LPIPS). LPIPS Zhang et al. [2018] is a modern perceptual similarity metric that is based on deep learning. It quantifies the similarity between two images by comparing their features at different layers of a neural network. This metric is believed to better correlate with human visual perception than traditional metrics by capturing perceptually important features. The LPIPS score between two images and is typically computed as:
@@
Structural Similarity Index Measure (SSIM). SSIM Wang and Bovik [2002] is a perceptual metric that assesses the visual impact of three characteristics: luminance, contrast, and structure. Instead of measuring pixel-wise differences, SSIM compares the structural information in the images, making it a more relevant measure of perceived quality. The SSIM index is calculated as follows:
@@
Appendix C Additional Experimental Results
C.1 Full Vbench Results.
VBench Huang et al. [2024b] is a comprehensive benchmark suite designed for evaluating video generative models. It provides a standardized framework that includes diverse evaluation metrics, datasets, and protocols to facilitate in-depth comparisons of model performance across various tasks. This systematic approach ensures reliable assessments of generated videos in terms of quality, efficiency, and robustness. Table 6 compares the performance of four video generation methods across each of the 16 VBench dimensions. A higher score indicates relatively better performance for a particular dimension. We also provide two specially built baselines, i.e., Empirical Min and Max (the approximated achievable min and max scores for each dimension), as references.
@@ Appendix C Additional Experimental Results
C.2 Last Step
| Method | Latency(s) | LPIPS | SSIM | PSNR |
| original, 19 steps | 29.20 | 0.3139 | 0.6922 | 16.39 |
| BWCache, 30 steps | 27.68 | 0.0879 | 0.8856 | 27.08 |
@@ C.2 Last Step
| Model | Method | Metrics | Mean | Standard Deviation | P-Value | Confidence Interval | ||
| Open-Sora | PAB | LPIPS | 0.2134 | 0.1026 | 9.90e-111 | [0.2048, 0.2220] | ||
| SSIM | 0.7798 | 0.1084 | 8.99e-73 | [0.7707, 0.7888] | ||||
| PSNR | 22.0191 | 2.8325 | 2.8e-135 | [21.7819, 22.2564] | ||||
| TeaCache | LPIPS | 0.1496 | 0.0780 | 1.07e-46 | [0.1430, 0.1561] | |||
| SSIM | 0.8104 | 0.1034 | 1.09e-42 | [0.8017, 0.8190] | ||||
| PSNR | 22.3929 | 4.3654 | 2.8e-80 | [22.0273, 22.7586] | ||||
| BWCache | LPIPS | 0.0879 | 0.0561 | - | [0.0832, 0.0926] | |||
| SSIM | 0.8854 | 0.0665 | - | [0.8798, 0.8909] | ||||
| PSNR | 27.0510 | 2.9898 | - | [26.8006, 27.3014] | ||||
PAB | LPIPS | 0.2781 | 0.1177 | 7.69e-51 | [0.2602, 0.2961] | ||
| SSIM | 0.6627 | 0.1449 | 7.16e-37 | [0.6407, 0.6848] | ||||
| PSNR | 19.4968 | 3.6949 | 6.76e-42 | [18.9340, 20.0596] | ||||
| TeaCache | LPIPS | 0.1965 | 0.0928 | 6.64e-27 | [0.1824, 0.2106] | |||
| SSIM | 0.7502 | 0.1090 | 4.95e-18 | [0.7336, 0.7668] | ||||
| PSNR | 21.5010 | 3.5785 | 6.25e-24 | [20.9559, 22.0461] | ||||
| BWCache | LPIPS | 0.1001 | 0.0520 | - | [0.0922, 0.1080] | |||
| SSIM | 0.8435 | 0.0743 | - | [0.8322, 0.8548] | ||||
| PSNR | 25.8734 | 3.7611 | - | [25.3006, 26.4463] | ||||
| Latte | PAB | LPIPS | 0.4625 | 0.1481 | 3.84e-218 | [0.4501, 0.4750] | ||
| SSIM | 0.5964 | 0.1631 | 3.88e-106 | [0.5827, 0.6100] | ||||
| PSNR | 17.0299 | 2.2196 | 2.67e-189 | [16.8440, 17.2158] | ||||
| TeaCache | LPIPS | 0.1969 | 0.1058 | 5.32e-17 | [0.1880, 0.2058] | |||
| SSIM | 0.7606 | 0.1329 | 2.39e-12 | [0.7495, 0.7718] | ||||
| PSNR | 22.1947 | 4.3783 | 1.02e-40 | [21.8280, 22.5614] | ||||
| BWCache | LPIPS | 0.1399 | 0.1160 | - | [0.1301, 0.1496] | |||
| SSIM | 0.8181 | 0.1357 | - | [0.8067, 0.8294] | ||||
| PSNR | 26.4620 | 5.6981 | - | [25.9847, 26.9392] |
C.3 Reducing Inference Steps
When generating videos using the DiT model, the impact of the inference steps must be considered. Although increasing the inference steps can enhance video quality, it clearly requires higher GPU usage and increases the generation latency. In this paper, the default setting for the inference steps in Open-Sora is 30. To accelerate video generation, the number of inference steps can be reduced. We compare the performance of the original model with reduced inference steps and BWCache in terms of visual quality and latency. The results show that, while producing similar latencies, BWCache maintains excellent visual quality, as shown in Figure 7. Specific data on visual quality and efficiency is detailed in Table 8. Overall, when the number of inference steps is reduced, sampling trajectories deviate from the theoretical optimal path, which amplifies accumulated errors in the noise estimation network. This results in issues such as blurred video details and reduced color saturation. BWCache maintains a high-fidelity video by caching and reusing intermediate stages of inference and their similar DiT block features, thus significantly reducing latency while accurately reconstructing the original inference process.
C.4 Statistical Analysis

### v3

@@ 3.4 Periodic Cache Recomputation
Continuously reusing the same block may result in latent drift and gradually erase fine-grained details Liu et al. (2025a). To prevent this cache stagnation, we adopt the progressive computation strategy from PAB Zhao et al. (2024b): within the caching interval, each DiT block is periodically recomputed at a defined reuse interval . Specifically, after computing the block at time step , its features are cached and reused for the subsequent steps (i.e., ), and the block is then updated at step . This process is formalized as:
@@ 4.1 Experiment Setup
Base Models and Compared Methods. We apply the proposed acceleration techniques to various video generation diffusion models, including Open-Sora Zheng et al. (2024), Open-Sora-Plan Lin et al. (2024a), Latte Ma et al. (2025a), Wan 2.1 Wan et al. (2025), and HunyuanVideo Kong et al. (2024). We compare our method with recent effective video generation techniques such as -DiT Chen et al. (2024b), T-GATE Zhang et al. (2024b), PAB Zhao et al. (2024b), and TeaCache Liu et al. (2024) to highlight our advantages. The inference configs of base models are shown in Appendix B.
Evaluation Metrics and Datasets. To evaluate the performance of video generation acceleration methods, we primarily focus on two aspects: visual quality and inference efficiency. We utilize VBench Huang et al. (2024b), LPIPS Zhang et al. (2018), PSNR, and SSIM Wang and Bovik (2002) for visual quality assessment. VBench is a comprehensive benchmarking suite for video generation models. LPIPS, PSNR, and SSIM measure the similarity between videos generated by the accelerated methods and the original models. Several evaluation metrics are detailed in Appendix C. To assess inference efficiency, we use the inference latency as metrics. In our main experiments, we generate over 1,000 videos for each model under each baseline using prompts sourced from Open-Sora gallery Lin et al. (2024a), VBench benchmark Huang et al. (2024b), and T2V-CompBench Sun et al. (2024).
@@ 4.2 Main Results
Quantitative Comparison. The experimental results in Table 1 demonstrate the effectiveness of BWCache across multiple DiT-based video generation models. Our proposed BWCache outperforms existing optimization methods in both visual quality and efficiency. For instance, on Open-Sora, BWCache achieves advanced visual metrics while reducing latency by 1.61 compared to the original model. Notably, BWCache surpasses the training-free TeaCache method and maintains competitive acceleration performance. The block-wise method of BWCache leverages DiT block redundancies more effectively without sacrificing visual details. The timestep-level caching of TeaCache is less adaptive to intra-timestep variations, leading to higher quality degradation. It is worth noting that TeaCache achieves faster acceleration than BWCache on the Open-Sora-Plan model. This is because there is significant variation in block features during the early inference stages of the Open-Sora-Plan model. To maintain high-fidelity output quality, BWCache strategically skips caching during the first few denoising steps where feature oscillations are most pronounced (as visualized in Figure 5(b)), thereby sacrificing some acceleration potential in favor of generation quality. We also compare with Skip-DiT Chen et al. (2025) on the Latte model, as shown in Table 1. While Skip-DiT achieves competitive acceleration (1.65 speedup), it requires architectural modifications and model retraining, which significantly limits its practical applicability. The Skip-DiT repository only provides a pre-trained model for Latte, and according to their documentation, training requires approximately 8 H100 GPUs for about one week with 300k text-video pairs. This substantial training requirement makes it impractical to evaluate Skip-DiT on other models or compare it fairly across different architectures. We further compare with FasterCache, which is another training-free caching-based acceleration method. The official FasterCache repository provides support for Open-Sora, Open-Sora-Plan, and Latte models. BWCache consistently outperforms FasterCache across all three models in terms of visual quality metrics. More experiments can be found in Appendix D.
@@ 4.3 Scaling Abilities
Scaling to Multiple GPUs. Table 3 summarizes the inference efficiency of various methods when scaled across multiple GPUs using Dynamic Sequence Parallelism (DSP) Zhao et al. (2024a). Utilizing the DSP method, GPUs are interconnected via high-bandwidth NVLink, which provides low-latency, high-throughput communication for distributed workloads. A single GPU often cannot utilize a full batch size due to memory constraints, leading to suboptimal hardware utilization. In contrast, the DSP method overcomes this by employing dimension-aware sharding and efficient all-to-all operations to minimize communication overhead, thereby enabling more effective scaling. The experiments, conducted using several models, reveal that BWCache consistently outperforms both the original model, PAB, and TeaCache in terms of latency reduction across all tested scenarios. Specifically, BWCache achieves the lowest latencies in almost all cases, demonstrating significant speed-ups as the number of GPUs increases. These results validate BWCache as a universally effective acceleration method for DiT-based models under multi-GPU deployments.
@@
| TeaCache | 17.5(1.5) | 7.3(3.5) | 5.6(4.66) | 5.0(5.2) |
| BWCache | 14.4(1.8) | 7.3(3.6) | 4.5(5.8) | 4.0(6.5) |
| Method | Latency(s) | LPIPS | SSIM | PSNR |
| original, 19 steps | 29.20 | 0.3139 | 0.6922 | 16.39 |
| BWCache, 30 steps | 27.68 | 0.0879 | 0.8856 | 27.08 |
Performance at Different Length and Resolution. Figure 7 demonstrates the inference efficiency of BWCache across various video lengths and resolutions using Open-Sora, Open-Sora-Plan, and Wan 2.1 models. The results show that latency increases significantly as the length and resolution of the video increase, but BWCache consistently performs better than the original model in all configurations. BWCache exhibits particularly notable acceleration advantages in processing high-resolution long videos. For example, with 8 GPUs, BWCache attains a latency of just 11.08 seconds for the Open-Sora model with 204 frames at 480P, representing a remarkable 17.16 speed-up compared to the base model. This trend is also reflected in other configurations, where BWCache consistently provides substantial improvements in efficiency. Overall, these results validate its capability to improve inference speeds across various video lengths and resolutions, making it an efficient solution for real-time video generation.
@@ 4.4 Ablation Studies
Reducing Inference Steps. When generating videos using the DiT model, the impact of the inference steps must be considered. In this paper, the default setting for the inference steps in Open-Sora is 30. To accelerate video generation, the number of inference steps can be reduced. We compare the performance of the original model with reduced inference steps and BWCache in terms of visual quality and latency. The results show that, while producing similar latencies, BWCache maintains excellent visual quality, as shown in Figure 8. Specific data on visual quality and efficiency is detailed in Table 3. Overall, when the number of inference steps is reduced, sampling trajectories deviate from the theoretical optimal path, which amplifies accumulated errors in the noise estimation network. This results in issues such as blurred video details and reduced color saturation.
Quality-Efficiency Trade-Off. Figure 1 compares the quality-latency trade-off of BWCache, PAB, and TeaCache. The indicator threshold in Eq.(7) is 0.15, 0.20, and 0.25. The results clearly demonstrate that BWCache outperforms PAB and TeaCache in visual quality across all evaluated metrics, indicating a notable improvement in the fidelity of video generation. For each metric, BWCache not only achieves higher quality scores but also maintains lower latency, which reflects its efficient processing capabilities. Overall, BWCache achieves improved visual quality at lower latency than PAB and TeaCache.
@@
Reuse Rate. Table 5 illustrates how varying the reuse rates affects performance. It reveals that the “Fast” configuration achieves the highest reuse rate at 59.00%, resulting in the lowest latency of 19.16 seconds. The “Slow” configuration achieves a reuse rate of 41.38%, but the visual quality is the highest among all configurations. Visualized results in Figure 9 complement these findings by providing the reuse rates of blocks across each timestep for various thresholds. This figure illustrates the dynamic nature of block reuse at different timesteps, with evident fluctuations that suggest varying block reuse efficiency at different stages of video generation. In general, the block features in the middle stage of the generation change less, and a high reuse rate is obtained, which is consistent with our previous analysis.
@@ 5 Conclusion
In this paper, we introduced BWCache, a novel caching method designed to accelerate video Diffusion Transformers. By analyzing feature dynamics in DiT blocks, we established that block features across timesteps had significant computational redundancy. BWCache addressed this by employing dynamic block-wise caching and a similarity indicator, which selectively reused block features when their differences fell below a predefined threshold. Periodic recomputation was employed to mitigate potential latent drift. Extensive experiments demonstrated that BWCache achieved robust efficiency and visual quality across diverse generation models, sampling schedules, and video parameters. These results highlighted its potential for real-world applications. The current indicator threshold was preset, and future work will dynamically adjust the threshold according to different generation tasks.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant 62272050 and Grant 62302048; in part by the Guangdong Key Lab of AI and Multi-modal Data Processing, Beijing Normal-Hong Kong Baptist University (BNBU), Zhuhai under 2023-2024 grants sponsored by Guangdong Provincial Department of Education; in part by the Institute of Artificial Intelligence and Future Networks (BNU-Zhuhai) and Engineering Center of AI and Future Education, Guangdong Provincial Department of Science and Technology, China; in part by Zhuhai Science-Tech Innovation Bureau under Grant No. 2320004002772; and in part by the Interdisciplinary Intelligent Supercomputing Center of Beijing Normal University (Zhuhai).
Appendix A Block-Level Feature Reuse
Figure 10 presents a comparative analysis of feature variations during the video generation process for the two distinct prompts introduced in Figure 4(4(a)). Figure 10(10(a)) illustrates the block-level feature differences, while Figure 10(10(b)) depicts the timestep-level feature differences. The results reveal that block-level features exhibit significant and discernible variations between the two prompts. Specifically, the video generated for the more static scene (e.g., Prompt 1) demonstrates lower feature variability, enabling more granular and aggressive caching and reuse of block features. This facilitates substantial acceleration while maintaining visual fidelity. In contrast, the timestep-level features show minimal differentiation across prompts, making it difficult to perform prompt-specific adaptive caching control. This fundamental limitation explains why timestep-level caching methods, such as TeaCache, often result in inferior video quality compared to our proposed block-wise method.
Appendix B Experimental Settings
Base Model. This paper focuses on DiT-based video generation, evaluating five state-of-the-art open-source models: Open-Sora Zheng et al. (2024), Open-Sora-Plan Lin et al. (2024a), Latte Ma et al. (2025a), Wan 2.1 Wan et al. (2025), and HunyuanVideo Kong et al. (2024). Open-Sora employs Spatial-Temporal DiT Ho et al. (2022), allowing for the creation of videos up to 16 seconds long at 720p resolution. Open-Sora Plan focuses on generating high-resolution videos with extended durations based on diverse user inputs, featuring components like a Wavelet-Flow Variational Autoencoder and specialized denoisers. Latte effectively extracts spatio-temporal tokens from videos and utilizes DiT blocks to model the token distribution in a latent space. Wan is a comprehensive and open-source suite of high-performance video foundation models. HunyuanVideo is an open-source video foundation model, integrating advanced data curation and a scalable transformer architecture. The inference configs of five models are shown in Table 6, which strictly follow the official settings.
@@ Appendix B Experimental Settings
| Wan 2.1 | FlowMatch | 50 | UMT5 |
| HunyuanVideo | FlowMatch | 50 | MLLM |
Appendix C Metrics
@@
Learned Perceptual Image Patch Similarity (LPIPS). LPIPS Zhang et al. (2018) is a modern perceptual similarity metric that is based on deep learning. It quantifies the similarity between two images by comparing their features at different layers of a neural network. This metric is believed to better correlate with human visual perception than traditional metrics by capturing perceptually important features. The LPIPS score between two images and is typically computed as:
@@
Structural Similarity Index Measure (SSIM). SSIM Wang and Bovik (2002) is a perceptual metric that assesses the visual impact of three characteristics: luminance, contrast, and structure. Instead of measuring pixel-wise differences, SSIM compares the structural information in the images, making it a more relevant measure of perceived quality. The SSIM index is calculated as follows:
@@
Appendix D Additional Experimental Results
D.1 Comparison with High-Memory Methods on a Single GPU.
We compare BWCache with high-memory caching methods, including ProfilingDiT Ma et al. (2025b) and TaylorSeer Liu et al. (2025b). These methods require caching extensive intermediate features, leading to significant memory overhead. When generating long videos, both ProfilingDiT and TaylorSeer encounter out-of-memory (OOM) errors. To ensure a fair comparison, we evaluate these methods on shorter video sequences where they can run successfully. Specifically, we test on Wan 2.1 with 9 frames at 480P and HunyuanVideo with 17 frames at 544P, which are the maximum video lengths that TaylorSeer can process on a single NVIDIA A800 80GB GPU, as shown in Table 7. We configure ProfilingDiT with , and TaylorSeer with and for Wan 2.1, and and for HunyuanVideo. Despite operating under more favorable conditions for the baseline methods, BWCache still achieves superior visual quality metrics while maintaining competitive or better acceleration performance. This demonstrates that BWCache provides a better balance between memory efficiency and generation quality compared to methods that require extensive feature caching.
| Model | Method | Visual Quality | Efficiency | ||||
| VBench | LPIPS | SSIM | PSNR | Speedup | Latency(s) | ||
| Wan 2.1 (9 frames, 480P) | Original | 77.05% | - | - | - | 1 | 287 |
| ProfilingDiT | 76.53% | 0.2053 | 0.8221 | 24.20 | 1.24 | 231 | |
| TaylorSeer | 75.13% | 0.3080 | 0.7231 | 18.75 | 1.87 | 153 | |
| BWCache | 76.81% | 0.0929 | 0.9048 | 29.65 | 1.67 | 171 | |
| HunyuanVideo (17 frames, 544P) | Original | 82.08% | - | - | - | 1 | 179 |
| ProfilingDiT | 80.33% | 0.2610 | 0.6662 | 18.77 | 1.72 | 104 | |
| TaylorSeer | 82.05% | 0.4021 | 0.5526 | 15.33 | 1.94 | 92 | |
| BWCache | 81.42% | 0.1804 | 0.7507 | 21.60 | 2.39 | 75 | |
D.2 Full Vbench Results.
VBench Huang et al. (2024b) is a comprehensive benchmark suite designed for evaluating video generative models. It provides a standardized framework that includes diverse evaluation metrics, datasets, and protocols to facilitate in-depth comparisons of model performance across various tasks. This systematic approach ensures reliable assessments of generated videos in terms of quality, efficiency, and robustness. Table 8 compares the performance of four video generation methods across each of the 16 VBench dimensions. A higher score indicates relatively better performance for a particular dimension. We also provide two specially built baselines, i.e., Empirical Min and Max (the approximated achievable min and max scores for each dimension), as references.
@@ Appendix D Additional Experimental Results
D.3 Last Step
@@
| Model | Method | Metrics | Mean | Standard Deviation | P-Value | Confidence Interval |
| Open-Sora | PAB | LPIPS | 0.2134 | 0.1026 | 9.90e-111 | [0.2048, 0.2220] |
| SSIM | 0.7798 | 0.1084 | 8.99e-73 | [0.7707, 0.7888] | ||
| PSNR | 22.0191 | 2.8325 | 2.8e-135 | [21.7819, 22.2564] | ||
| TeaCache | LPIPS | 0.1496 | 0.0780 | 1.07e-46 | [0.1430, 0.1561] | |
| SSIM | 0.8104 | 0.1034 | 1.09e-42 | [0.8017, 0.8190] | ||
| PSNR | 22.3929 | 4.3654 | 2.8e-80 | [22.0273, 22.7586] | ||
| BWCache | LPIPS | 0.0879 | 0.0561 | - | [0.0832, 0.0926] | |
| SSIM | 0.8854 | 0.0665 | - | [0.8798, 0.8909] | ||
| PSNR | 27.0510 | 2.9898 | - | [26.8006, 27.3014] | ||
| Open-Sora- Plan | PAB | LPIPS | 0.2781 | 0.1177 | 7.69e-51 | [0.2602, 0.2961] |
| SSIM | 0.6627 | 0.1449 | 7.16e-37 | [0.6407, 0.6848] | ||
| PSNR | 19.4968 | 3.6949 | 6.76e-42 | [18.9340, 20.0596] | ||
| TeaCache | LPIPS | 0.1965 | 0.0928 | 6.64e-27 | [0.1824, 0.2106] | |
| SSIM | 0.7502 | 0.1090 | 4.95e-18 | [0.7336, 0.7668] | ||
| PSNR | 21.5010 | 3.5785 | 6.25e-24 | [20.9559, 22.0461] | ||
| BWCache | LPIPS | 0.1001 | 0.0520 | - | [0.0922, 0.1080] | |
| SSIM | 0.8435 | 0.0743 | - | [0.8322, 0.8548] | ||
| PSNR | 25.8734 | 3.7611 | - | [25.3006, 26.4463] | ||
| Latte | PAB | LPIPS | 0.4625 | 0.1481 | 3.84e-218 | [0.4501, 0.4750] |
| SSIM | 0.5964 | 0.1631 | 3.88e-106 | [0.5827, 0.6100] | ||
| PSNR | 17.0299 | 2.2196 | 2.67e-189 | [16.8440, 17.2158] | ||
| TeaCache | LPIPS | 0.1969 | 0.1058 | 5.32e-17 | [0.1880, 0.2058] | |
| SSIM | 0.7606 | 0.1329 | 2.39e-12 | [0.7495, 0.7718] | ||
| PSNR | 22.1947 | 4.3783 | 1.02e-40 | [21.8280, 22.5614] | ||
| BWCache | LPIPS | 0.1399 | 0.1160 | - | [0.1301, 0.1496] | |
| SSIM | 0.8181 | 0.1357 | - | [0.8067, 0.8294] | ||
| PSNR | 26.4620 | 5.6981 | - | [25.9847, 26.9392] | ||
| Wan 2.1 | PAB | LPIPS | 0.1623 | 0.1141 | 2.10e-18 | [0.1440, 0.1806] |
| SSIM | 0.7643 | 0.1465 | 4.92e-09 | [0.7405, 0.7881] | ||
| PSNR | 21.3210 | 3.5476 | 3.07e-22 | [20.7974, 21.8446] | ||
| TeaCache | LPIPS | 0.2382 | 0.0894 | 1.49e-50 | [0.2232, 0.2531] | |
| SSIM | 0.6608 | 0.1406 | 1.47e-31 | [0.6373, 0.6843] | ||
| PSNR | 18.6197 | 3.2942 | 1.11e-42 | [18.0692, 19.1702] | ||
| BWCache | LPIPS | 0.0805 | 0.0461 | - | [0.0728, 0.0882] | |
| SSIM | 0.8527 | 0.0968 | - | [0.8365, 0.8689] | ||
| PSNR | 25.7246 | 3.9350 | - | [25.0670, 26.3821] | ||
| HunyuanVideo | PAB | LPIPS | 0.1045 | 0.1224 | 3.20e-06 | [0.0857, 0.1233] |
| SSIM | 0.8341 | 0.1642 | 1.47e-03 | [0.8070, 0.8613] | ||
| PSNR | 26.6958 | 3.6582 | 2.11e-11 | [26.1089, 27.2827] | ||
| TeaCache | LPIPS | 0.1685 | 0.0823 | 3.68e-25 | [0.1547, 0.1822] | |
| SSIM | 0.7973 | 0.1184 | 2.80e-13 | [0.7775, 0.8171] | ||
| PSNR | 23.9004 | 3.9613 | 1.14e-25 | [23.2384, 24.5623] | ||
| BWCache | LPIPS | 0.0768 | 0.0468 | - | [0.0689, 0.0846] | |
| SSIM | 0.8922 | 0.0860 | - | [0.8778, 0.9065] | ||
| PSNR | 29.6333 | 4.2949 | - | [28.9157, 30.3510] |
D.4 Statistical Analysis